OpenStudy (anonymous):

Let u = <-3, -5>, v = <-3, 1>. Find u + v. answers: <-2, -8> <-8, -2> <0, -6> <-6, -4>

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

@lanybug96 do you want to learn math?

OpenStudy (anonymous):

I have no idea, super lost

OpenStudy (skullpatrol):

$$\Huge u+v$$ $$\Huge =<x_u,y_u>+<x_v,y_v>$$ $$\Huge=<x_u+x_v,y_u+y_v>$$

OpenStudy (skullpatrol):

Add the x-components and add the y-components.

OpenStudy (skullpatrol):

$$\Huge u = <-3, -5>, v = <-3, 1>$$

OpenStudy (skullpatrol):

-3 +(-3) = ? -5 + 1 = ?

OpenStudy (skullpatrol):

Do you want me to go over the rules for addition using negative numbers?

OpenStudy (skullpatrol):

If two numbers have the same sign add their absolute values and prefix their common sign.

OpenStudy (anonymous):

okay is it B? sorry! my computer kept doing weird things and I couldn't reply

OpenStudy (skullpatrol):

nope, not <-8, -2>

OpenStudy (anonymous):

okay well when I add -3 + -5 I get -8 and -3 + 1 I get -2... so whats the stepp after that?

OpenStudy (skullpatrol):

Each vector has two components, right?

OpenStudy (anonymous):

right

OpenStudy (skullpatrol):

The first component is called the x-component The second is called the y-component

OpenStudy (skullpatrol):

$$\Huge<x,y>$$

OpenStudy (anonymous):

$$\Huge<x,y>$$ what is that suppose to mean?

OpenStudy (skullpatrol):

x is first number and y is the second number

OpenStudy (skullpatrol):

just like a variable

OpenStudy (skullpatrol):

a symbol used to represent a number

OpenStudy (skullpatrol):

$$\Huge <1st,2nd>$$

OpenStudy (anonymous):

okay lol I get that part... im just lost5 on how to find answer

OpenStudy (anonymous):

is it D?

OpenStudy (skullpatrol):

yes